OpenStudy (anonymous):

There are six boys and six girls at a party. In how many ways can they be seated in a row if they must sit alternating boys and girls?

OpenStudy (anonymous):

Total number of outcomes is 12!. Possibilities for boys' seating is 6! and possibilities for girls' seating is 6!. 2*(6!) = 2*720 = 1440